# Vibrationally Assisted Exciton Transfer in Open Quantum Systems with Long-Range Interactions

### v1

@@ III.2 Non-Perturbative regime
Figure 3: Optimal transfer and static disorder: (a) The transfer rate as a function of $J_{TT}/\gamma$ is presented for different parameter conditions. $J$ is kept fixed for each plot and $\gamma$ is varied. The reference parameter set (REF in the legend) corresponds to $J=0.3\omega$, $p=1$, $g=\omega$, $\epsilon=3\omega$, and $\bar{n}=0.01$. For all other curves, one of these parameters is modified, as indicated in the legend, while all others remain the same. Four $\gamma$ values are chosen for further analysis (see Table 2 for the exact values of $\gamma$). These parameter sets are labeled as $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$, corresponding to the square, circle, diamond, and triangle markers, respectively. The horizontal dashed line is used to illustrate that $p_{2}$, $p_{3}$, and $p_{4}$ yield almost identical transfer rates. (b) Transfer rate as a function of the standard deviation of the static disorder in the $g_{j}$ couplings. $k_{T}*$ represents the transfer value without disorder. Each data point is the average over one hundred disorder realizations, and error bars represent the 25th and 75th percentiles. (c) Same as in (b) but for disordered on-site energies $\epsilon_{j}$.
III.3 Optimal Transfer and Static Disorder
@@
To study the effect of static disorder on the transfer rate, we consider four parameter configurations $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$, signaled with markers in Fig. 3(a), and we average over 100 disorder realizations. $p_{1}$ represents the optimal $J_{TT}/\gamma$ ratio, while $p_{2}$, $p_{3}$, and $p_{4}$ are chosen such that they all yield almost identical transfer rates despite representing different parameter values (see horizontal dashed line in Fig. 3(a)). To model the static disorder, we consider $g_{j}=\bar{g}_{j}(1+\delta^{g}_{j})$, with $\bar{g}_{1}=\bar{g}_{2}=-\bar{g}_{3}=-\bar{g}_{4}=g$, and $\delta^{g}_{j}$ being a Gaussian random variable with zero average and standard deviation $\sigma_{g}$. In Fig. 3(b), results for different values of $\sigma_{g}$ are shown. The vertical axis is normalized with respect to the optimal value without disorder $k_{T}^{*}$ for each parameter set $p_{j}$. As expected, when the disorder increases, the transfer rate is lowered. A site-dependent $g_{j}$ changes the energy of the vibronic states (see Eq. (8)), affecting the resonance condition. Furthermore, the disorder in the displacements $g_{j}/\omega$ associated with the donor (acceptor) sites also reduces the transfer rate by decreasing the Frank-Condon factor. The decrease in transfer rate can be understood by noting that the reorganization energies of each qubit $\lambda_{j}=g_{j}^{2}/\omega$ are, on average, always larger than the electronic coupling $J$ for our parameter choice, resulting in dynamical localization of the excitonic states.
As $\sigma_{g}$ increases, we note a higher variance of the possible transfer rates. Indeed, in the presence of static disorder, the different $g_{j}$ can favor transfer processes that are not resonant when all couplings have the same magnitude $|g_{j}|=g$. This may result in the prominence of slower transfer triplet-to-singlet or singlet-to-singlet processes. Moreover, certain disorder realizations can lead to a larger donor population in the steady state, decreasing the transfer rate (see discussion after Eq. 14 and Appendix C). Additionally, the transfer rates in the four configurations in Fig. 3(b) are all affected in a similar way by disorder in the spin phonon-coupling $g_{j}$. However, the optimal parameter set $p_{1}$, on average, always yields a higher transfer rate in absolute terms since the vertical axis in Fig. 3(b) is normalized with respect to $k_{T}^{*}$.
@@
A difference in $\epsilon_{j}$ only modifies the energy shifts of the different states, differently from the site-dependent $g_{j}$ couplings that also modify the Franck-Condon factors. In general, we expect to still have delocalized states contributing to the dynamics if the energy variations are small compared to the electronic coupling ($\sigma_{\epsilon}\ll J/\omega$), while a more disordered energy landscape ($\sigma_{\epsilon}\gg J/\omega$) localizes the excitations, decreasing the transfer rate. Additionally, as shown in Fig. 2(c), a faster relaxation rate $\gamma$ broadens the width of the resonances in the spectrum, therefore making the resonance conditions more robust against variations in $\epsilon_{j}$. The results in Fig. 3(c) confirm this behavior as the parameter set $p_{3}$ (circle markers) with $\gamma\approx 0.1\omega$ exhibits a considerably slower decrease in the transfer rate than all other parameter choices with $0.01\lesssim\gamma/\omega\lesssim 0.04$. Interestingly, at high disorder $\sigma_{\epsilon}=0.1$, the parameter set $p_{3}$ with larger $\gamma$ exhibits a larger transfer $k_{T}$ also in absolute terms compared to the other configurations, differently from the case of small disorder, when $p_{1}$ configuration is optimal. Consequently, for fixed electronic coupling $J_{TT}$, the presence of static disorder changes the relaxation rate $\gamma$ that maximizes the transfer. The interplay of dissipation and disorder has also been explored in the case of Anderson localization [62] and light-harvesting complexes [63].
@@ IV Effect of temperature
where $\bar{n}$ corresponds to the temperature set by the bath and all the calculations are performed with $j=T_{D}$. We report in Fig. 4(a) the transfer rate as a function of $\epsilon$ for different temperatures. First, when the temperature increases, the transfer rate at the resonant peaks (integer values of $\epsilon/\omega$) decreases. However, it is important to point out that the transfer rates at higher values of $\epsilon$ are more resilient to the temperature increase than those for lower $\epsilon$.
We can understand this behavior through the schematics presented in Fig. 4(b). Initially, a thermal state (Eq. (16)) is coherently coupled to a family of states $\{|\psi_{T_{A},n}\rangle\}$ with coupling strength $J_{TT}$. In the acceptor sites, the dissipative dynamics couple different vibrational states until a thermal distribution is reached in the steady state. Indeed, the heating term in the Eq. (2) is relevant when $\bar{n}$ is close to 1.
@@ V.1 Larger Monomers
where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are normalization constants (see Appendix D). Similarly, acceptor states $|\mathcal{E}^{1}_{A}\rangle$, $|\mathcal{E}^{2}_{A}\rangle$, and $|\mathcal{E}^{3}_{A}\rangle$ can be equivalently expressed in terms of the states $|j_{A}\rangle$. We can write the Hamiltonian in Eq. (1) in this basis (see Appendix D for details). The resulting couplings can be written in the notation $J_{AB}$ where $A,B=\mathcal{E}^{1},\mathcal{E}^{2},\mathcal{E}^{3}$.
As shown explicitly in Appendix D, the larger of the inter-monomer couplings for $p=1$ is $J_{\mathcal{E}^{1}\mathcal{E}^{1}}$. Importantly, the state $|\mathcal{E}^{1}_{D}\rangle$ has a large overlap with the $W$ state $|W_{D}\rangle=\frac{1}{\sqrt{3}}(|1\rangle+|2\rangle+|3\rangle)$. Similar to the $L=2$ case studied previously, the transfer is also favored by an initial symmetric superposition of product states if $J_{ij}>0$. Since the $W$ state is a paradigmatic state in quantum information science [69] that can be efficiently prepared with digital protocols [70], we study how transfer from an initial $|W_{D}\rangle$ state behaves in comparison to the initial state $|\mathcal{E}^{1}_{D}\rangle$. Specifically, in Fig. 5(a-b), we study how the transfer rate changes as the number of qubits per monomer $L$ is increased. The number of eigenstates $|\mathcal{E}^{j_{D}}_{D}\rangle$ increases linearly with the number of qubits in each monomer $L$. Here, for consistency, $|\mathcal{E}^{1}_{D}\rangle$ is defined as the state with the largest overlap to the corresponding $W$ state for each $L$. As shown in Appendix D, this state has the highest eigenenergy and always yields the largest inter-monomer coupling strength for any monomer size $L$ considered here.
In Fig. 5(a), we show the dependence of the transfer rate defined in Eq. (14) on the number of qubits per monomer $L$. As expected, for larger system sizes, the transfer rate decreases. For each $L$, the transfer from the initial state $|\mathcal{E}^{1}_{D}\rangle$ is always comparable to the $W$ state because of the large overlap of these two states. However, since the fastest processes favor the transfer to the $|\mathcal{E}_{A}^{1}\rangle$ state (see Appendix D), the final state of the system will not be the $|W_{A}\rangle$ state, but it will be closer to $|\mathcal{E}_{A}^{1}\rangle$, even when the system is initialized in $|W_{D}\rangle$. This is illustrated in the dynamics depicted in Fig. 5(b).
@@ VI.1 Trapped-Ion Model of Excitation Transfer
Individual Raman beams with multiple frequency beatnotes of different amplitudes and phases on the qubit ions can be used to generate the desired Hamiltonian in Eq. (1). A frequency beatnote on resonance with the qubit frequency $\omega_{0}$ (carrier drive) allows for individual energy shifts to modulate the energy landscape of the system, $\sum_{i}\sigma_{i}^{z}\epsilon_{i}/2$, in the appropriate spin basis. The site-to-site electronic coupling term, $\sum_{i<j}J_{ij}(\sigma_{i}^{+}\sigma_{j}^{-}+\sigma_{i}^{-}\sigma_{j}^{+})$, can be realized natively by simultaneously driving two symmetric beatnotes $\omega_{0}\pm(\omega_{N}+\delta_{J})$, where $\omega_{N}$ is the secular frequency of the center-of-mass collective mode. This results in a long-range Ising XX interaction with the coupling strength $J_{ij}=J/d_{ij}^{p}$, where $p$ can be adjusted by changing the axial frequency of the ion chain and the detuning $\delta_{J}$. A strong transverse field $B\sum_{i}\sigma^{z}_{i}$ ($J\ll B\ll\delta_{J}$) can be applied to suppress the double excitation processes present in the native Ising interaction, therefore enforcing the conservation of the total number of spin excitations along the transverse field direction. This can be realized either by shifting the two symmetric sideband frequency tones used for generating the Ising interaction with a common detuning or by implementing an additional qubit frequency resonant tone, depending on the spin basis chosen for the Hamiltonian engineering [49]. In the following section, which describes an experimental setup for simulating the two-monomer case, we demonstrate the use of the latter approach. The residual $B$ field only adds a global offset to the energy landscape of the system and does not affect its dynamics.
@@
Finally, it shall also be noted that the proposed scheme to generate the vibronic coupling and the harmonic terms can off-resonantly couple to other collective motional modes $(m\neq k)$ and induce additional two-body interactions. However, these terms result only in state-dependent shifts to the energy landscape of the system and do not affect its single electronic excitation manifold description. These energy shifts can be corrected experimentally by adjusting the single-site energy offsets (see Appendix E).
@@ VI.2 Two-monomer model
As an alternative strategy to scale up to larger systems, it is possible to use qudits instead of qubits, where we encode the monomer’s electronic configuration with more than two internal states of the ions [78]. However, this comes at the cost of engineering complex electronic interactions with additional lasers to create the desired Hamiltonian. Lastly, a key component of the simulation of more complex monomers proposed here is the preparation of $W$ states (or states with a large overlap with $W$ states), which is achievable through digital [70] or analog methods [81]. This hybrid digital-analog approach is also ideal for implementing the recently proposed simulation of 2D spectroscopy using a probe qubit [82], which significantly reduces measurement overhead compared to standard digital protocols [83].

### v2

@@ III.2 Non-Perturbative regime
Figure 3: Optimal transfer and static disorder: (a) The transfer rate as a function of $J_{TT}/\gamma$ is presented for different parameter conditions. $J$ is kept fixed for each plot and $\gamma$ is varied. The reference parameter set (REF in the legend) corresponds to $J=0.3\omega$, $p=1$, $g=\omega$, $\epsilon=3\omega$, and $\bar{n}=0.01$. For all other curves, one of these parameters is modified, as indicated in the legend, while all others remain the same. Four $\gamma$ values are chosen for further analysis (see Table 2 for the exact values of $\gamma$). These parameter sets are labeled as $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$, corresponding to the square, circle, diamond, and triangle markers, respectively. The horizontal dashed line is used to illustrate that $p_{2}$, $p_{3}$, and $p_{4}$ yield almost identical transfer rates. (b) Transfer rate as a function of the standard deviation of the static disorder in the $g_{j}$ couplings. $k_{T}*$ represents the transfer value without disorder. Each data point is the average over one hundred disorder realizations, and error bars represent the 25th and 75th percentiles. (c) Same as in (b) but for disordered on-site energies $\epsilon_{j}$. (d) Transfer rate as a function of the dephasing rate $\gamma_{d}$. $k_{T}*$ represents the transfer rate value without noise. All markers and colors are defined as in (b) and (c).
III.3 Optimal Transfer, Static Disorder, and Noise
@@
To study the effect of static disorder on the transfer rate, we consider four parameter configurations $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$, signaled with markers in Fig. 3(a), and we average over 100 disorder realizations. $p_{1}$ represents the optimal $J_{TT}/\gamma$ ratio, while $p_{2}$, $p_{3}$, and $p_{4}$ are chosen such that they all yield almost identical transfer rates despite representing different parameter values (see horizontal dashed line in Fig. 3(a)). To model the static disorder, we consider $g_{j}=\bar{g}_{j}(1+\delta^{g}_{j})$, with $\bar{g}_{1}=\bar{g}_{2}=-\bar{g}_{3}=-\bar{g}_{4}=g$, and $\delta^{g}_{j}$ being a Gaussian random variable with zero average and standard deviation $\sigma_{g}$. In Fig. 3(b), results for different values of $\sigma_{g}$ are shown. The vertical axis is normalized with respect to the optimal value without disorder $k_{T}^{*}$ for each parameter set $p_{j}$. As expected, when the disorder increases, the transfer rate is lowered. A site-dependent $g_{j}$ changes the energy of the vibronic states (see Eq. (8)), affecting the resonance condition. Furthermore, the disorder in the displacements $g_{j}/\omega$ associated with the donor (acceptor) sites also reduces the transfer rate by decreasing the Franck-Condon factor. The decrease in transfer rate can be understood by noting that the reorganization energies of each qubit $\lambda_{j}=g_{j}^{2}/\omega$ are, on average, always larger than the electronic coupling $J$ for our parameter choice, resulting in dynamical localization of the excitonic states.
As $\sigma_{g}$ increases, we note a higher variance of the possible transfer rates. Indeed, in the presence of static disorder, the different $g_{j}$ can favor transfer processes that are not resonant when all couplings have the same magnitude $|g_{j}|=g$. This may result in the prominence of slower transfer triplet-to-singlet or singlet-to-singlet processes. Moreover, certain disorder realizations can lead to a larger donor population in the steady state, decreasing the transfer rate (see discussion after Eq. 14 and Appendix E). Additionally, the transfer rates in the four configurations in Fig. 3(b) are all affected in a similar way by disorder in the spin-phonon coupling $g_{j}$. However, the optimal parameter set $p_{1}$, on average, always yields a higher transfer rate in absolute terms since the vertical axis in Fig. 3(b) is normalized with respect to $k_{T}^{*}$.
@@
A difference in $\epsilon_{j}$ only modifies the energy shifts of the different states, differently from the site-dependent $g_{j}$ couplings that also modify the Franck-Condon factors. In general, we expect to still have delocalized states contributing to the dynamics if the energy variations are small compared to the electronic coupling ($\sigma_{\epsilon}\ll J/\omega$), while a more disordered energy landscape ($\sigma_{\epsilon}\gg J/\omega$) localizes the excitations, decreasing the transfer rate. Additionally, as shown in Fig. 2(c), a faster relaxation rate $\gamma$ broadens the width of the resonances in the spectrum, therefore making the resonance conditions more robust against variations in $\epsilon_{j}$. The results in Fig. 3(c) confirm this behavior as the parameter set $p_{3}$ (circle markers) with $\gamma\approx 0.1\omega$ exhibits a considerably slower decrease in the transfer rate than all other parameter choices with $0.01\lesssim\gamma/\omega\lesssim 0.04$. Interestingly, at high disorder $\sigma_{\epsilon}=0.1$, the parameter set $p_{3}$ with larger $\gamma$ exhibits a larger transfer $k_{T}$ also in absolute terms compared to the other configurations, differently from the case of small disorder, when the $p_{1}$ configuration is optimal. Consequently, for fixed electronic coupling $J_{TT}$, the presence of static disorder changes the relaxation rate $\gamma$ that maximizes the transfer. The interplay of dissipation and disorder has also been explored in the case of Anderson localization [68] and light-harvesting complexes [69].
Apart from static disorder, time-dependent noise in the system parameters has shown to be important in different transfer setups [70, 71, 72], especially in the context of environment-assisted quantum transfer (ENAQT), where moderate noise strengths might enhance excitation transfer rates [73, 74, 69]. In our case, we incorporate noise by considering the parameters $\epsilon_{j}(t)$ to change in time. As shown in Ref. [75], if the fluctuations on $\epsilon_{j}(t)$ can be described by white-noise stochastic processes, the dynamics can be equivalently modeled by a time-independent Hamiltonian and a set of Lindbladian jump operators of the form $\mathcal{L}_{\sigma_{j}^{z}}[\rho]$ (see Eq. (3)). Here we consider the same approach and add the terms $\gamma_{d}\sum_{j=1}^{N}\mathcal{L}_{\sigma_{j}^{z}}[\rho]$ to the master equation in Eq. (2), where the dephasing rate $\gamma_{d}$ is proportional to the width of the white noise. Here, we consider the same dephasing rate for all qubits.
In Fig. 3(d) we show the normalized transfer rate as a function of $\gamma_{d}$ for the four sets of parameters $p_{j}$ in Table 2. In general, dephasing leads to a slower transfer rate, as compared to the noise-free case. However, the dependence on $\gamma_{d}$ is not monotonic for all cases. For parameter sets $p_{1}$, $p_{2}$, and $p_{3}$, the transfer rapidly reaches a minimum, and then it increases again up to a local maximum. Similar cases in which the transfer rate has a non-monotonic behavior as a function of the noise strength have been reported in similar setups [70, 72]. We should note, however, that we do not observe a case for the chosen parameter sets, where the presence of noise leads to a faster transfer with respect to the noise-free case. This is expected as the transfer speed in the model studied here is enhanced due to the coherences between the qubits of each monomer, and dephasing naturally destroys these coherences.
@@ IV Effect of temperature
where $\bar{n}$ corresponds to the temperature set by the bath, and all the calculations are performed for the triplet donor state with $|\psi_{j,n}\rangle=|\psi_{T_{D},n}\rangle$. We report in Fig. 4(a) the transfer rate as a function of $\epsilon$ for different temperatures. First, when the temperature increases, the transfer rate at the resonant peaks (integer values of $\epsilon/\omega$) decreases. However, it is important to point out that the transfer rates at higher values of $\epsilon$ are more resilient to the temperature increase than those for lower $\epsilon$.
We can understand this behavior through the schematics presented in Fig. 4(b). Initially, a thermal state (Eq. (16)) is coherently coupled to a family of states $\{|\psi_{T_{A},n}\rangle\}$ with coupling strength $J_{TT}$. In the acceptor sites, the dissipative dynamics couples different vibrational states until a thermal distribution is reached in the steady state. Indeed, the heating term in Eq. (2) is relevant when $\bar{n}$ is close to 1.
@@ V.1 Larger Monomers
where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are normalization constants (see Appendix F). Similarly, acceptor states $|\mathcal{E}^{1}_{A}\rangle$, $|\mathcal{E}^{2}_{A}\rangle$, and $|\mathcal{E}^{3}_{A}\rangle$ can be equivalently expressed in terms of the states $|j_{A}\rangle$. We can write the Hamiltonian in Eq. (1) in this basis (see Appendix F for details). The resulting couplings can be written in the notation $J_{AB}$ where $A,B=\mathcal{E}^{1},\mathcal{E}^{2},\mathcal{E}^{3}$.
As shown explicitly in Appendix F, the larger of the inter-monomer couplings for $p=1$ is $J_{\mathcal{E}^{1}\mathcal{E}^{1}}$. Importantly, the state $|\mathcal{E}^{1}_{D}\rangle$ has a large overlap with the $W$ state $|W_{D}\rangle=\frac{1}{\sqrt{3}}(|1\rangle+|2\rangle+|3\rangle)$. Similar to the $L=2$ case studied previously, the transfer is also favored by an initial symmetric superposition of product states if $J_{ij}>0$. Since the $W$ state is a paradigmatic state in quantum information science [81] that can be efficiently prepared with digital protocols [82], we study how transfer from an initial $|W_{D}\rangle$ state behaves in comparison to the initial state $|\mathcal{E}^{1}_{D}\rangle$. Specifically, in Fig. 5(a-b), we study how the transfer rate changes as the number of qubits per monomer $L$ is increased. The number of eigenstates $|\mathcal{E}^{j_{D}}_{D}\rangle$ increases linearly with the number of qubits in each monomer $L$. Here, for consistency, $|\mathcal{E}^{1}_{D}\rangle$ is defined as the state with the largest overlap to the corresponding $W$ state for each $L$. As shown in Appendix F, this state has the highest eigenenergy and always yields the largest inter-monomer coupling strength for any monomer size $L$ considered here.
In Fig. 5(a), we show the dependence of the transfer rate defined in Eq. (14) on the number of qubits per monomer $L$. As expected, for larger system sizes, the transfer rate decreases. For each $L$, the transfer from the initial state $|\mathcal{E}^{1}_{D}\rangle$ is always comparable to the $W$ state because of the large overlap of these two states. However, since the fastest processes favor the transfer to the $|\mathcal{E}_{A}^{1}\rangle$ state (see Appendix F), the final state of the system will not be the $|W_{A}\rangle$ state, but it will be closer to $|\mathcal{E}_{A}^{1}\rangle$, even when the system is initialized in $|W_{D}\rangle$, as is illustrated in the dynamics depicted in Fig. 5(b). Given that the system is initialized in the entangled state $|W_{A}\rangle$ and evolves almost completely into the state $|\mathcal{E}_{A}^{1}\rangle$, a large portion of the entanglement is transferred from donor to acceptor, similar to the case of simpler monomers (see Appendix D for more details).
@@ VI.1 Trapped-Ion Model of Excitation Transfer
Individual Raman beams with multiple frequency beatnotes of different amplitudes and phases on the qubit ions can be used to generate the desired Hamiltonian in Eq. (1). A frequency beatnote on resonance with the qubit frequency $\omega_{0}$ (carrier drive) allows for individual energy shifts to modulate the energy landscape of the system, $\sum_{i}\sigma_{i}^{z}\epsilon_{i}/2$, in the appropriate spin basis. The site-to-site electronic coupling term, $\sum_{i<j}J_{ij}(\sigma_{i}^{+}\sigma_{j}^{-}+\sigma_{i}^{-}\sigma_{j}^{+})$, can be realized natively by simultaneously driving two symmetric beatnotes $\omega_{0}\pm(\omega_{N}+\delta_{J})$, where $\omega_{N}$ is the secular frequency of the center-of-mass collective mode. This results in a long-range Ising XX interaction with the coupling strength $J_{ij}=J/d_{ij}^{p}$, where $p$ can be adjusted by changing the axial frequency of the ion chain and the detuning $\delta_{J}$. The tunable range of $p$ is between 0 and 3 [49], but the experimentally accessible range is approximately between 0.5 and 2. A strong transverse field $B\sum_{i}\sigma^{z}_{i}$ ($J\ll B\ll\delta_{J}$) can be applied to suppress the double excitation processes present in the native Ising interaction, therefore enforcing the conservation of the total number of spin excitations along the transverse field direction. This can be realized either by shifting the two symmetric sideband frequency tones used for generating the Ising interaction with a common detuning or by implementing an additional qubit frequency resonant tone, depending on the spin basis chosen for the Hamiltonian engineering [49]. In the following section, which describes an experimental setup for simulating the two-monomer case, we demonstrate the use of the latter approach. The residual $B$ field only adds a global offset to the energy landscape of the system and does not affect its dynamics.
@@
We also note that the proposed scheme to generate the vibronic coupling and the harmonic terms can off-resonantly couple to other collective motional modes $(m\neq k)$ and induce additional two-body interactions. However, these terms result only in state-dependent shifts to the energy landscape of the system and do not affect its single electronic excitation manifold description. These energy shifts can be corrected experimentally by adjusting the single-site energy offsets (see Appendix G).
Finally, the simultaneous application of the four sideband tones in both bichromatic drives can induce A.C. Stark shifts, which introduce spurious terms into the engineered trapped-ion Hamiltonian, deviating away from the desired model. Therefore, in practice, one also must compensate for these shifts with a proper choice of the amplitudes and frequencies of the laser tones, which have been successfully achieved by several analog trapped-ion experiments with a similar setup using two simultaneous bichromatic tones [39, 66, 90].
@@ VI.2 Two-monomer model
As an alternative strategy to scale up to larger systems, it is possible to use qudits instead of qubits, where we encode the monomer’s electronic configuration with more than two internal states of the ions [91]. However, this comes at the cost of engineering complex electronic interactions with additional lasers to create the desired Hamiltonian.
Lastly, a key component of the simulation of more complex monomers proposed here is the preparation of $W$ states (or states with a large overlap with $W$ states), which is achievable through digital [82] or analog methods [94]. This hybrid digital-analog approach is also ideal for implementing the recently proposed simulation of 2D spectroscopy using a probe qubit [95], which significantly reduces measurement overhead compared to standard digital protocols [96].